# A fully Bayesian approach for the imputation and analysis of derived outcome variables with missingness

### v1

@@ Abstract
Derived variables are variables that are constructed from one or more source variables through established mathematical operations or algorithms. For example, body mass index (BMI) is a derived variable constructed from two source variables: weight and height. When using a derived variable as the outcome in a statistical model, complications arise when some of the source variables have missing values. In this paper, we propose how one can define a single fully Bayesian model to simultaneously impute missing values and sample from the posterior. We compare our proposed method with alternative approaches that rely on multiple imputation, and, with a simulated dataset, consider how best to estimate the risk of microcephaly in newborns exposed to the ZIKA virus.
@@ 1 Introduction
The data that motivated this work consists of measurements taken from newborns of mothers infected with the ZIKA virus (Wilder-Smith et al. 2019). In order to determine if the newborns have microcephaly (a derived variable), a calculation based on three source variables (sex, gestational age, and head circumference) is required. The source variables are often subject to missingness and we were curious as to how one might be able to fit these data with an entirely Bayesian solution.
Besides the so-called complete-case analysis, in which observations with any missing values are discarded, the most popular statistical approach for handling missing data is multiple imputation (MI) (Rubin 1976, Reiter & Raghunathan 2007) and among the different varieties of MI, the Multivariate Imputation by Chained Equations (MICE) approach is the most widely used (Van Buuren 2018). The basic idea of MI is to fill in missing values by repeated simulation from the posterior predictive distributions of the missing values given assumptions about the missing data mechanism and a model, thereby preserving the uncertainty associated with the missingness. MICE does this for multivariate non-monotone missing data variables by using a sequence of univariate conditional distributions (i.e., imputing variables one-at-a-time).
When using MICE to impute missing values of derived variables, there are three general approaches:
@@
    imputation at the source-variable level (SVL), in which one imputes the source variables first and then computes the derived variable afterwards (also known as the impute, then transform approach); and
@@
    the just another variable (JAV), approach, in which one imputes the source variables and derived variable together, with the source variables as covariates in the imputation model for the derived variable.
There are two other approaches we should mention that have been studied in the literature but are not necessarily relevant for our situation. The first is known as “on-the-fly” imputation, where the transformation of the source variables into the derived variable is done “on-the-fly” within the imputation algorithm. (Note that Van Buuren (2018) calls this “passive imputation”). This can be advantageous when there are observations for which the derived variable is observed but the source variables are missing (e.g., there are individuals with known BMI, but unknown height and/or weight). We will assume that missingness in source variables necessarily implies missingness in the derived variable (e.g., any individual with unknown height or weight will necessarily have unknown BMI). In this case, on-the-fly imputation is entirely equivalent to SVL imputation. The other approach we should mention, proposed by Bartlett et al. (2015), is known as the Substantive Model Compatible Fully Conditional Specification (SCMFCS) imputation. This method is useful for imputing derived covariate values, but not derived outcome variables (which is our focus).
Because DVL imputation effectively excludes information available in the source variables (which are possibly the best predictors of the missing derived variable values), SVL imputation is thought to be a more efficient strategy, in the sense that it uses observed information where possible. Pan et al. (2020) compare the performance of the two approaches and conclude that SVL imputation is indeed preferable, regardless of how the derived variable is obtained (“whether by simple arithmetic operations or by some highly specific algorithms”). This follows related work by Gottschall et al. (2012) who arrived at a similar conclusion for handling missing values in questionnaire data (“researchers should adopt item-level imputation whenever possible”).
@@
Another reason to favour the SVL and JAV approaches over DVL imputation is that they make a less restrictive MAR (missing at random) assumption. MAR requires that the conditional probability of missingness depends only on data observed under that pattern (Rubin 1976). Consider for example a derived outcome $Y$ that is equal to the sum of two source variables, $Z_{1}$ and $Z_{2}$, such that $Y=Z_{1}+Z_{2}$. Missingness in $Y$ can occur in three patterns: First, the $Z_{1}$ value is missing, second the $Z_{2}$ value is missing, third, both $Z_{1}$ and $Z_{2}$ values are missing; see Table 1. Because MAR is i) defined at the observation level rather than at the variable level, and ii) conditional on variables included in the analysis, we see that assuming MAR with DVL imputation requires that missingness in the derived variable $Y$ be independent of the value of $Z_{1},Z_{2}$ under any pattern. For the SVL and JAV approaches, this is relaxed under patterns 1 and 2: missingness in $Z_{1}$ under pattern 1 need only be independent of $Z_{1}$, not $Z_{2}$, and missingness in $Z_{2}$ under pattern 2 need only be independent of $Z_{2}$, not $Z_{1}$.
@@
To conduct Bayesian inference with imputed data, Gelman et al. (2004) outline a three step procedure. The first step is to construct several imputed datasets, using MI. The second step is to simulate draws from the posterior distribution with each imputed dataset separately. The third and final step is to mix all the draws together, thereby obtaining (at least approximately) draws from the posterior distribution that take into account the uncertainty associated with the missing data; see Zhou & Reiter (2010). (This third step is perhaps surprising to those familiar with multiple imputation inference: Rubin’s combining rules are not needed.) While Gelman et al. (2004)’s three step procedure appears straightforward, when SVL imputation is used for imputing a derived outcome variable, things can be much more complicated. This becomes apparent when considering the assumptions required in the first two steps.
@@
Using SVL imputation in the first step requires one to specify distributions for each of the missing source variables. Then, specifying the posterior distribution in second step requires one to define a distribution – along with priors for the parameters that characterize the distribution – for the derived outcome variable. Clearly the assumed distributions of the source variables (specified in step 1) necessarily determine the distribution of the derived outcome variable (which must be specified in step 2). Without careful consideration, one could inadvertently define incompatible distributions, since the two steps are done separately and the software implementing one step (e.g., MICE from van Buuren et al. (2015)) does not know of the assumptions required by the other step (e.g., rstan from Carpenter et al. (2017)). Moreover, it will no doubt be difficult to determine whether or not the assumptions made by the Bayesian model specified in the second step (e.g., regarding priors) are consistent with the assumptions made when imputing missing values in the first step.
@@
Ideally, one could fit a single fully Bayesian model to simultaneously impute missing values and sample from the posterior. The idea of avoiding potential problems with MI by adopting a fully Bayesian approach is not new and it is actually quite common to proceed in this way when covariate values are missing; see Bartlett et al. (2015) and Erler et al. (2021) and the references therein. However, it is not at all obvious how to proceed when there is missingness in source variables that combine to define a derived outcome variable. In this article, we propose a useful approach.
@@
In Section 2, we begin with a simple illustrative example. In Section 3, we outline our proposed approach. In Section 4, we consider how best to estimate the risk of microcephaly in newborns exposed to the ZIKA virus with an artificial dataset. Finally, we conclude in Section 5.
2 Illustrative example
@@
We continue with the simple example discussed in the introduction where $Y=Z_{1}+Z_{2}$. Suppose we are interested in two groups, group $A$ and group $B$, (e.g., placebo and treatment) and we are interested in estimating $\theta$, the difference between $\theta_{A}$, the expectation of $Y$ in group $A$, and $\theta_{B}$, the expectation of $Y$ in group $B$: $\theta=\theta_{A}-\theta_{B}$. The easiest way to proceed might be to fit a simple univariate Normal model for $Y$ where, for $i=1,\ldots,N$:
$y_{i}$ $\sim\text{Normal}(\theta_{i},\sigma_{y}^{2})\quad$
where
$\boldsymbol{\theta_{i}}$ =$=\begin{cases}\theta_{A}&\text{if observation $i$ is in group $A$},\\ \theta_{B}&\text{if observation $i$ is in group $B$},\\ \end{cases}$
with priors specified for $\theta_{A}$, $\theta_{B}$ and $\sigma_{y}$.
@@
Now suppose that, in group $A$, a proportion of $Z_{1}$ and $Z_{2}$ values are missing. Specifically, in group $A$, certain observations are prone to have missing $Z_{1}$ values with the chance of missingness depending on $Z_{2}$, and other observations are prone to have missing $Z_{2}$ values with the chance of missingness depending on the observed $Z_{1}$. (To be clear, in this specific example, no observations are susceptible to missing both $Z_{1}$ and $Z_{2}$.) Table 3 lists the data for six individuals to illustrate, with “NA” indicating missing. Given this missingness in the data, how should we proceed to estimate $\theta$?
$i$ Group $Y$ $Z_{1}$ $Z_{2}$
@@
As discussed in the introduction, one could consider the complete-case analysis in which case, all rows of data with any missing values would be discarded (e.g., individuals $i=1$ and $i=3$ listed in Table 3 would be discarded). The advantage of this strategy is that no imputation step is required and the entire procedure is to simply fit the univariate Normal model. The disadvantage is that this strategy almost certainly involves a loss of efficiency and has the potential for bias.
Alternatively, one could apply Gelman’s three step procedure with one of the three MICE approaches (either the DVL, SVL, or JAV imputation methods) for imputing the missing $Y$ values followed by the univariate Normal model for inference. The advantage of this strategy is that, depending on the specific MICE approach used, one could conceivably obtain unbiased estimates of $\theta$ in an efficient manner. The disadvantage is that this strategy involves two separate steps (an imputation step and an inference step) which may not be consistent with one another.
There is another possible strategy worth considering. Because the math in this scenario is rather simple, one could fit a Bayesian model for the source variables, $Z_{1}$ and $Z_{2}$, and derive samples of $\theta$ based on samples of $Z_{1}$ and $Z_{2}$. For instance, we could fit a bivariate Normal model for $Z_{1}$ and $Z_{2}$, where, for $i$ in 1,…,$N$:
@@
This strategy might be ideal in the sense that it does not require a separate imputation step (the entire procedure is to simply fit the bivariate Normal model) and does not disregard a large number of observations. However, we this strategy is possible only because in this illustrative example the math is extremely straightforward (i.e., equation (2) is known). For other derived variables, this might not be the case. For instance, if instead of $Y=Z_{1}+Z_{2}$, we had $Y=Z_{1}/({Z_{2}}^{2})$, then computing $\theta$ from a combination of the seven other parameters might be impossible without a certain degree of mathematical wizardry.
This brings us to our proposed method, details of which are provided in the next section. Briefly, the approach involves using Monte Carlo integration to derive samples of the derived outcome variable. In our example this would involve fitting the bivariate Normal model and sampling $\theta$ (at least approximately) without requiring any knowledge about the mathematical derivation of $\theta$.
Before continuing on with more details, let us review the results of a small simulation study to confirm that our intuitions regarding each of the possible strategies in the illustrative example are acceptable. We simulated 500 datasets of size $N=1,000$ and applied each of the six methods (details are in the Appendix). The aim was to demonstrate a proof-of-concept for our proposed method (A7). Five hundred complete-data observations were generated as
$\left(\begin{array}[]{c}Z_{1}\\ Z_{2}\end{array}\right)$ $\sim\textrm{Normal}\begin{pmatrix}\left(\begin{array}[]{c}1\\ 2\end{array}\right),\left(\begin{array}[]{cc}1&0.25\\ 0.25&1\end{array}\right)\end{pmatrix},$ if $i$ is in group $A$
$\left(\begin{array}[]{c}Z_{1}\\ Z_{2}\end{array}\right)$ $\sim\textrm{Normal}\begin{pmatrix}\left(\begin{array}[]{c}1.5\\ 1\end{array}\right),\left(\begin{array}[]{cc}1&0.25\\ 0.25&1\end{array}\right)\end{pmatrix},$ if i is in group $\text{if $i$ is in group $B$}.$
Subsequently, missingness was imposed under a missing-at-random mechanism. All data remained fully observed for group $B$. For group $A$, 250 observations had non-zero probability of $z_{1}$ being missing and the other 250 had non-zero probability of $z_{2}$ being missing. The following models determined the probability of missingness in these two subsets.
$\text{logit}[P(R_{1}=0)]$ $=10+10z_{2},$
and  $\text{and~{}~{}logit}[P(R_{2}=0)]$ $=4-5z_{1}.$
For analysis of the incomplete data, each of the MICE methods was used to produce 50 imputed datasets, each Bayesian model was fit using MCMC with $M=5,000$ draws, and the proposed method was implemented with $S=50,000$ Monte Carlo samples for the integration. We also analysed data prior to introduction of missingness (i.e., from the oracle) and fit the univariate normal model to these data. Figure 1 shows the results. In summary, we see that:
  • The complete-case analysis (A2) and the three-step approach with DVL imputation (A4) are both biased as expected, since the data generating mechanism is not MCAR.
  • The three-step approach with SVL imputation (A3) is unbiased.
  • Both approaches that fit the bivariate Normal model (A6 and A7) are unbiased and just as efficient as the three-step approach with SVL imputation (A3).
  • Finally, the three-step approach with JAV (A5) imputation appears to be biased. This is surprising, since in this case JAV imputation should be identical to SVL imputation. A brief explanation of this claim is given in the Appendix. In short, this is not a property of JAV imputation but of the R package {mice} used to implement it.
Figure 1: Simulation results - We simulated 500 datasets of size $N=1,000$ with true $\theta=-0.5$ and applied each of the six methods (A2–A7) recording the posterior mean estimate. We also analyzed data before introducing missingness (i.e., from the oracle) and fit the univariate normal model to these data. For reference: A1. = Oracle with univariate Normal model; A2. = Complete case analysis with univariate Normal model; A3. = Gelman’s 3 step with SVL imputation and univariate Normal model; A4. = Gelman’s 3 step with DVL imputation with univariate Normal model; A5. = Gelman’s 3 step with JAV imputation with univariate Normal model; A6. = Bivariate Normal model (with math); and A7. = Bivariate Normal model (with MC approx.).
@@ 3 Description of the approach
Having established the motivation behind our proposed method, we now outline the details more specifically. Note that, while we outline the method as it applies to an analysis in which the objective is to determine the relationship between a derived outcome variable and a single binary exposure variable, the method could be modified as needed to suit other scenarios.
For the $i$-th observation in $i$=1,…,$N$, let $X_{i}$ be the exposure variable, $Y_{i}$ be the derived outcome variable, and $Z_{i,1},...,Z_{i,P}$ be $P$ source variables, such that $Y_{i}=f(Z_{i,1},...,Z_{i,P})$, for some deterministic function $f()$. For instance, in the earlier example, $X_{i}$ was the “group” membership and the function $f()$ defined the addition of $P=2$ source variables: $f(Z_{i,1},Z_{i,2})=Z_{i,1}+Z_{i,2}$.
Suppose $\theta$ is the estimand of interest and the primary goal is in determining the posterior distribution of $\theta$. In the earlier example, interest was in determining the difference in the mean of $Y$ between individuals in the two groups, such that $\theta=\textrm{E}(Y|X=1)-\textrm{E}(Y|X=0)$.
Now suppose that there is missingness for various values of $Z_{1},\ldots,Z_{P}$, and that that we know (or can safely assume) the conditional distribution for the source variables, $(Z|X)$. This is the key assumption on which our proposal rests. Let $\phi$ be all the parameters describing the $(Z|X)$ distribution and let $h()$ be a function (perhaps unknown) such that $\theta=h(\phi)$. In the previous example, a bivariate Normal model was specified for $(Z|X)$ (see equation (1)), $\phi$ consisted of seven parameters ($\mu_{z1,A}$, $\mu_{z2,A}$, $\mu_{z1,B}$, $\mu_{z2,B}$ $\sigma_{z1}$, $\sigma_{z2}$ and $\rho$), and $h()$ was known and defined as: $h(\phi)=(\mu_{z1,A}+\mu_{z2,A})-(\mu_{z1,B}+\mu_{z2,B})$ (see equation (2)).
The standard Bayesian approach would be to obtain a Monte Carlo sample of size $M$ (e.g., via MCMC) from the $(\phi,Z_{mis}|Z_{obs},X)$ posterior distribution and then, if $h()$ is known, obtain $M$ posterior samples of $\theta$ by applying the $h()$ function to the $M$ posterior samples of $\phi$. (In the previous section, this was the A6 “Bivariate Normal model (with math)” approach). Instead, our proposal is to create an additional variable, $Z^{*}_{p}$, for each of the $P$ source variables. For $p$ in 1,…,$P$, we define $Z^{*}_{p}$ such that its distribution is identical to that of $Z_{p}$. However, whereas $Z_{p}$ is partially (or entirely) observed, $Z^{*}_{p}$ is entirely unobserved. In this sense, $Z^{*}_{p}$ could be understood as “future observations projected by the model” of $Z_{p}$. With a sufficiently large number $S$ of draws of $Z^{*}_{p}$, the distribution of $Z^{*}_{p}$ should accurately approximate that of $Z_{p}$.
@@
By then applying the deterministic function $f()$ on samples of $Z^{*}_{1},...,Z^{*}_{P}$, one can obtain samples of $Y^{*}$, a new variable with the same identical distribution as $Y$ (i.e., “future observations projected by the model” of the $Y$ variable). Then, one can obtain posterior samples of $\theta$ based on the $Y^{*}$ values. Obtaining samples of $\theta$ from $Y^{*}$ in this fashion ensures that the correct amount of variation is captured and, importantly, does not require knowledge of the function $h()$. To be clear, our approach consists of the following four steps:
@@
    Draw $M$ samples of $\phi$ from the $(\phi|Z_{obs},X)$ distribution: $\phi^{[m]}$, for $m$ in 1,…,$M$.

### v2

@@ Abstract
Derived variables are variables that are constructed from one or more source variables through established mathematical operations or algorithms. For example, body mass index (BMI) is a derived variable constructed from two source variables: weight and height. When using a derived variable as the outcome in a statistical model, complications arise when some of the source variables have missing values. In this paper, we propose how one can define a single fully Bayesian model to simultaneously impute missing values and sample from the posterior. We compare our proposed method with alternative approaches that rely on multiple imputation with examples including an analysis to estimate the risk of microcephaly (a derived variable based on sex, gestational age and head circumference at birth) in newborns exposed to the ZIKA virus.
@@ 1 Introduction
The data that initially motivated this work consists of measurements taken from newborns of mothers infected with the ZIKA virus (Wilder-Smith et al. 2019). In order to determine if the newborns have microcephaly (a derived variable), a calculation based on three source variables (sex, gestational age, and head circumference) is required. The source variables are often subject to missingness and we were curious as to how one might be able to fit these data with an entirely Bayesian solution.
Besides so-called complete-case analysis, in which observations with any missing values are discarded, the most popular statistical approach for handling missing data is multiple imputation (MI) (Rubin 1976, Reiter & Raghunathan 2007) and among the different varieties of MI, the Multivariate Imputation by Chained Equations (MICE) approach is the most widely used (Van Buuren 2018). The basic idea of MI is to fill in missing values by repeated simulation from the posterior predictive distributions of the missing values given assumptions about the missing data mechanism and a model, thereby preserving the uncertainty associated with the missingness. MICE does this for multivariate non-monotone missing data variables by using a sequence of univariate conditional distributions (i.e., imputing variables one-at-a-time).
When using MICE to impute missing values of derived variables, there are four general approaches:
@@
    imputation at the source-variable level (SVL), in which one imputes the source variables first and then computes the derived variable afterwards (also known as the impute, then transform approach);
@@
    the just another variable (JAV), approach, in which one imputes the source variables and derived variable together, with the source variables as covariates in the imputation model for the derived variable. To be clear, all variables are imputed separately, ignoring any deterministic relationships. Instead, a multivariate normality between the derived variable and the sources variables is often assumed (e.g., Von Hippel (2009) and White et al. (2011)). Notably, Seaman et al. (2012) conclude that JAV should not be used to impute data for a logistic regression analysis as it performs “very badly” in this case; and
  4. 4.
    the “on-the-fly” imputation, where the transformation of the source variables into the derived variable is done “on-the-fly” within the imputation algorithm. (Note that Van Buuren (2018) calls this “passive imputation”). The “on-the-fly” imputation consists of three steps. First, the source variables are imputed, then the derived variable is updated based on the known deterministic relationship with the source variables, and, third, all other variables are imputed conditional on this updated derived variable. When missingness only occurs in the derived variable and source variables (i.e., no other variables need imputing) and missingness in source variables necessarily implies missingness in the derived variable (e.g., any individual with unknown height or weight will necessarily have unknown BMI), on-the-fly imputation is entirely equivalent to SVL imputation.
The other approach we should mention, proposed by Bartlett et al. (2015), is known as the Substantive Model Compatible Fully Conditional Specification (SCMFCS) imputation. This method is useful for imputing derived covariate values, but not derived outcome variables (which is our focus).
There may be challenges when imputing at the source-variable level if the derived variable is not a linear function of the source variables; see Seaman et al. (2012), Morris et al. (2014), and Tilling et al. (2016). However, because DVL imputation effectively excludes information available in the source variables (which are possibly the best predictors of the missing derived variable values), SVL (or “on-the-fly”) imputation is thought to be a more efficient strategy, in the sense that it uses observed information where possible. Pan et al. (2020) compare the performance of the DVL and SVL approaches and conclude that SVL imputation is indeed preferable, regardless of how the derived variable is obtained (“whether by simple arithmetic operations or by some highly specific algorithms”). This follows related work by Gottschall et al. (2012) who arrived at a similar conclusion for handling missing values in questionnaire data (“researchers should adopt item-level imputation whenever possible”).
@@
Another reason to favour the SVL, “on-the-fly”, and JAV approaches over DVL imputation is that they make a less restrictive MAR (missing at random) assumption. MAR requires that the conditional probability of missingness depends only on data observed under that pattern (Rubin 1976). Consider for example a derived outcome $Y$ that is equal to the sum of two source variables, $Z_{1}$ and $Z_{2}$, such that $Y=Z_{1}+Z_{2}$. Missingness in $Y$ can occur in three patterns: First, the $Z_{1}$ value is missing, second the $Z_{2}$ value is missing, third, both $Z_{1}$ and $Z_{2}$ values are missing; see Table 1. Because MAR is i) defined at the observation level rather than at the variable level, and ii) conditional on variables included in the analysis, we see that assuming MAR with DVL imputation requires that missingness in the derived variable $Y$ be independent of the value of $Z_{1},Z_{2}$ under any pattern, see Table 2. For the SVL, “on-the-fly”, and JAV approaches, this is relaxed under patterns 1 and 2: missingness in $Z_{1}$ under pattern 1 need only be independent of $Z_{1}$, not $Z_{2}$, and missingness in $Z_{2}$ under pattern 2 need only be independent of $Z_{2}$, not $Z_{1}$.
$A=Z_{1}$ (source) $A=Z_{2}$ (source) $A=Y$ (derived)
@@
To conduct Bayesian inference with imputed data, Gelman et al. (2004) outline a three step procedure. The first step is to construct $K$ imputed datasets, using MI. The second step is to simulate draws from the posterior distribution of the model parameters with each imputed dataset separately. The third step is to mix all the draws together, thereby obtaining (at least approximately) draws from the posterior distribution that take into account the uncertainty associated with the missing data; see Zhou & Reiter (2010). Summaries (e.g., median, 2.5th, and 97.5th percentiles) of the mix of draws from across the imputed datasets are then obtained for one’s parameter estimates. (This third step is perhaps surprising to those familiar with multiple imputation inference: Rubin’s combining rules are not needed.) While Gelman et al. (2004)’s three-step procedure appears straightforward, when SVL (or “on-the-fly”) imputation is used for imputing a derived outcome variable, things can be much more complicated. This becomes apparent when considering the assumptions required in the first two steps.
@@
Using SVL or “on-the-fly” imputation in the first step requires one to specify distributions for each of the missing source variables. Then, specifying the posterior distribution in second step requires one to define a distribution – along with priors for the parameters that characterize the distribution – for the derived outcome variable. Clearly the assumed distributions of the source variables (specified in step 1) necessarily determine the distribution of the derived outcome variable (which must be specified in step 2). Without careful consideration, one could inadvertently define incompatible distributions, since the two steps are done separately and the software implementing one step (e.g., MICE from van Buuren et al. (2015)) does not know of the assumptions required by the other step (e.g., rstan from Carpenter et al. (2017)). Moreover, it will no doubt be difficult to determine whether or not the assumptions made by the Bayesian model specified in the second step (e.g., regarding priors) are consistent with the assumptions made when imputing missing values in the first step.
@@
Ideally, one could fit a single fully Bayesian model to simultaneously impute missing values and sample from the posterior. The idea of avoiding potential problems with MI by adopting a fully Bayesian approach is not new and it is actually quite common to proceed in this way when covariate values are missing; see Bartlett et al. (2015), Lüdtke et al. (2020) and Erler et al. (2021) and the references therein. However, it is not at all obvious how to proceed when there is missingness in source variables that combine to define a derived outcome variable. In this article, we propose a useful approach.
@@
In Section 2, we begin with a simple illustrative example. In Section 3, we outline our proposed approach. In Section 4, we consider two applied examples: (1) comparing the BMI of Dutch boys from inside the city to those from outside the city, and (2) estimating the risk of microcephaly in newborns exposed to the ZIKA virus with an artificial dataset. Finally, we conclude in Section 5.
2 A simple illustrative example
@@
We continue with the simple example discussed in the introduction where $Y=Z_{1}+Z_{2}$. Suppose we are interested in two groups, group $A$ and group $B$, (e.g., placebo and treatment) and we are interested in estimating the difference between the mean of $Y$ in group $A$, and the mean of $Y$ in group $B$. The easiest way to proceed might be to fit a simple univariate Normal model for $Y$ where, for $i=1,\ldots,n$:
$y_{i}$ $\sim\text{Normal}(\alpha+\beta x_{i},\sigma_{y}^{2})\quad$ (1)
where $x_{i}=0$ if observation $i$ is in group $A$, and $x_{i}=1$ if observation $i$ is in group $B$, with priors specified for $\alpha$, $\beta$ and $\sigma_{y}$. The estimand (the target quantity that is to be estimated) coincides with $\beta$.
@@
Now suppose that, in group $A$, a proportion of $Z_{1}$ and $Z_{2}$ values are missing. Specifically, in group $A$, certain observations are prone to have missing $Z_{1}$ values with the chance of missingness depending on $Z_{2}$, and other observations are prone to have missing $Z_{2}$ values with the chance of missingness depending on the observed $Z_{1}$. (To be clear, in this specific example, no observations are susceptible to missing both $Z_{1}$ and $Z_{2}$.) Table 3 lists the data for six individuals to illustrate, with “NA” indicating missing. Given this missingness in the data, how should we proceed to estimate $\beta$?
$i$ X (Group) $Y$ $Z_{1}$ $Z_{2}$
@@
As discussed in the introduction, one could consider the complete-case analysis, where all rows of data with any missing values are discarded (individuals $i=1$ and $i=3$ listed in Table 3). The advantage of this strategy is that no imputation step is required and the entire procedure is to simply fit the univariate Normal model. The disadvantage is that this strategy almost certainly involves a loss of efficiency and has the potential for bias.
Alternatively, one could apply Gelman’s three step procedure with one of the three MICE approaches (either the DVL, SVL, or JAV imputation methods) for imputing the missing $Y$ values followed by the univariate Normal model for inference. The advantage of this strategy is that, depending on the specific MICE approach used, one could conceivably obtain unbiased estimates of $\beta$ in an efficient manner. The disadvantage is that this strategy involves two separate steps (an imputation step and an inference step) which may not be consistent with one another.
There is another possible strategy worth considering. Because the math in this scenario is rather simple, one could fit a Bayesian model for the source variables, $Z_{1}$ and $Z_{2}$, and derive the “implied” samples of $\beta$. For instance, we could fit a bivariate Normal model for $Z_{1}$ and $Z_{2}$, where, for $i$ in 1,…,$n$:
@@
This strategy might be ideal in the sense that it does not require a separate imputation step (the entire procedure is to simply fit the bivariate Normal model) and does not disregard a large number of observations. However, it is possible only because in this simple illustrative example the math is extremely straightforward (i.e., equation (3) is known). For other derived variables, this might not be the case. For instance, if instead of $Y=Z_{1}+Z_{2}$, we had $Y=Z_{1}/({Z_{2}}^{2})$, then computing $\beta$ from a combination of the seven model parameters ($\mu_{Z1,A}$, $\mu_{Z2,A}$, $\mu_{Z1,B}$, $\mu_{Z2,B}$ $\sigma_{Z1}$, $\sigma_{Z2}$ and $\rho$) might be impossible without a degree of mathematical wizardry.
This brings us to our proposed method, details of which are provided in the next section. Briefly, the approach involves using Monte Carlo integration to derive samples of the derived outcome variable. In our example this would involve fitting the bivariate Normal model and sampling $\beta$ (at least approximately) without requiring any knowledge about the mathematical derivation of $\beta$.
@@ 3 Description of the approach
Having established the motivation behind our proposed method, we now outline the details.
Suppose the data consists of $n$ observations and let $W$ be the set of $p+q$ variables at hand. Let $Y=f(W)=f(W_{source})$ be the outcome variable, where $W_{source}$ are the $p$ source variables that combine in some deterministic way to equal the outcome variable. (To be clear, there are $p$ source variables which make up the outcome $Y$). Let $W_{expl}$ consist of the remaining $q$ explanatory variables. In practice, note that certain variables could be both source variables and explanatory, but we set this aside to avoid more complicated notation. Here, $\dim(W_{source})=(n,p)$ and $\dim(W_{expl})=(n,q)$.
Say $W$ partitions as variables $W_{1}$ which are prone to missingness, and variables $W_{2}$ which are not: $W=\{W_{1},W_{2}\}$. Specifically, $W_{1}$ consists of certain observations that are missing, $W_{1}^{\{mis\}}$, and certain observations that are observed, $W_{1}^{\{obs\}}$, such that $W_{1}=\{W_{1}^{\{mis\}},W_{1}^{\{obs\}}\}$. We are interested in the situation in which there is overlap between $W_{1}$ and $W_{source}$, i.e., situations when certain source variables are subject to missingness. Let $W_{1,source}=W_{1}\cap W_{source}$, $W_{1,expl}=W_{1}\cap W_{expl}$, $W_{2,source}=W_{2}\cap W_{source}$, and $W_{2,expl}=W_{2}\cap W_{expl}$.
In the simple illustrative example we have $W=\{Z_{1},Z_{2},X\}$ (where $X$ is the binary group indicator variable such that, for the $i$-th observation, $X_{i}=0$ if $i$-th observation is in Group A and $X_{i}=1$ if the $i$-th observation is in Group B). Furthermore, we have $W_{1}=\{Z_{1},Z_{2}\}$, $W_{2}=\{X\}$, $p=2$, $W_{source}=\{Z_{1},Z_{2}\}$, $q=1$, $W_{expl}=\{X\}$, and $Y=f(W)=Z_{1}+Z_{2}$.
Typically, the objective of a Bayesian analysis consists of the estimation of an estimand (Kahan et al. 2023). The estimand of interest, $\theta$, is often defined non-parametrically as a function of certain conditional expectations: $\theta=g(\textrm{E}^{*}_{1}\{\textrm{E}(Y|W_{expl})\},...,\textrm{E}^{*}_{D}\{\textrm{E}(Y|W_{expl})\})$, where $\textrm{E}^{*}_{d}$ corresponds to the counterfactual expectation with respect to the $d$-th population of interest for $d$ in 1,…,$D$.
In certain situations, the target population(s) might themselves be unknown and require estimation. For example, suppose the estimand of interest is the mean difference in $Y$ between those receiving treatment ($X=1$) and those not receiving treatment ($X=0$) for individuals who are of “average age”, where average age, $\mu_{A}$, is unknown. Then, we have $D=2$ target populations: (1) treated individuals of average age, and (2) untreated individuals of average age. The estimand is defined as: $\theta=\textrm{E}(Y|X=1,A=\mu_{A})-\textrm{E}(Y|X=0,A=\mu_{A})$. The parameter $\mu_{A}$ is required to define the two target populations and must be estimated from the data.
In the simple illustrative example, the estimand of interest is the difference in the mean of $Y$ between individuals in the two groups (i.e., the difference between the conditional expectation of $Y$ given the $D=2$ target populations defined by $X=0$ and $X=1$): $\theta=\textrm{E}(Y|X=1)-\textrm{E}(Y|X=0)$. Because the math is conveniently simple, one can also define this estimand analytically as a function of the model parameters: $\theta=\beta$, but also as: $\theta=h(\phi_{source})=(\mu_{Z1,A}+\mu_{Z2,A})-(\mu_{Z1,B}+\mu_{Z2,B})$, where $\phi_{source}=\{\mu_{Z1,A}$, $\mu_{Z2,A}$, $\mu_{Z1,B}$, $\mu_{Z2,B}$ $\sigma_{Z1}$, $\sigma_{Z2}$, $\rho\}$. The standard Bayesian approach would then be to obtain a Monte Carlo sample of size $M$ (e.g., via MCMC) from the $(\phi_{source},W_{1}^{\{mis\}}|W_{1}^{\{obs\}},W_{2})$ posterior distribution and then, since $h()$ is known, obtain $M$ posterior samples of $\theta$ by applying the $h()$ function to each of the $M$ posterior samples of $\phi_{source}$.
@@
In situations when the math is not so simple, Bayesian g-computation (Gustafson 2015, Keil et al. 2018) can be used for a form of ‘model-based standardization’ (Zou 2009, Vansteelandt & Keiding 2011) in order to obtain posterior samples of $\theta$ as follows. First, for each of the target populations of interest, one forward samples “posterior predictive values” of the source variables conditional on the target population of interest. Then, by applying the deterministic function $f()$ on the sample of these “posterior predictive values” one can obtain a sample of $Y^{*}_{d}$, a new variable with the same identical distribution as $Y$ under the $d$-th population of interest, i.e., a “future observation” projected by the model of the $Y$ variable from the $d$-th target population. Finally, with a large number, $S$, of these “future observations” one can approximate $\textrm{E}^{*}_{d}(\textrm{E}(Y|W_{expl}))$, for $d$ in 1,…$D$, and combine these according to the definition of the estimand of interest to obtain a posterior sample of $\theta$. Obtaining samples of $\theta$ in this way ensures that the correct amount of variation is captured and, importantly, does not require knowledge of a function $h:\phi_{source}\rightarrow\theta$.
To be clear this approach consists of the following steps:
@@
    Draw $M$ samples of $(\phi_{source},W_{1,expl}^{\{mis\}})$ from the $(\phi_{source},W_{1,expl}^{\{mis\}}|W_{1}^{\{obs\}},W_{2})$ distribution: $(\phi_{source},W_{1,expl}^{\{mis\}})^{[m]}$, for $m$ in 1,…,$M$.